# compat=1.1.0

### v1

@@ IV Large neutrino asymmetry and $N_{\rm eff}$
In the SM, such neutrino asymmetries do not arise and hence there is no additional contribution to $N_{\rm eff}$. The sole contribution to $N_{\rm eff}$ in the SM comes from relativistic nature of light neutrinos. If neutrinos decouple instantaneously, then $N_{\rm eff}=3$. Considering non-instantaneous decoupling of SM neutrinos along with flavour oscillations and plasma correction of quantum electrodynamics, the SM $N_{\rm eff}$ value shifts to $N_{\rm eff}^{\rm SM}=3.045$ Mangano:2005cc ; Grohs:2015tfy ; deSalas:2016ztq . A deviation from $N_{\rm eff}^{\rm SM}$ indicates presence of BSM physics either in terms of additional light degrees of freedom, new interactions of SM neutrinos or large neutrino asymmetry mentioned above. The current bound on $N_{\rm eff}$ from Planck 2018 data is given as $N_{\rm eff}=2.99^{+0.34}_{-0.33}$ at $2\sigma$ CL including baryon acoustic oscillation (BAO) data. This corresponds to $\Delta N_{\rm eff}\lesssim 0.285$. The latest DESI 2024 data give a slightly weaker bound $\Delta N_{\rm eff}\lesssim 0.4$ at $2\sigma$ CL DESI:2024mwx . Similar bound also exists from BBN considerations $3.42.3<{\rm N}_{\rm eff}<$ 3.4 at $95\%$ CL Cyburt:2015mya . Future CMB experiment CMB Stage IV (CMB-S4) is expected reach a much better sensitivity of $\Delta{\rm N}_{\rm eff}={\rm N}_{\rm eff}-{\rm N}^{\rm SM}_{\rm eff}=0.06$ Abazajian:2019eic , taking it closer to the SM prediction. Another future experiment CMB-HD CMB-HD:2022bsz can probe $\Delta N_{\rm eff}$ upto $0.014$ at $1\sigma$.
In addition to enhancing $N_{\rm eff}$, a large neutrino asymmetry can also affect the BBN predictions of light nuclei abundance, as pointed out recently in the light of anomalous observations related to primordial Helium-4 (${}^{4}{\rm He}$) abundance. The recent observations made by the Subaru Survey Matsumoto:2022tlr , together with previous observations, have indicated a shift in primordial ${}^{4}{\rm He}$ abundance $Y_{P}=0.2379^{+0.0031}_{-0.0030}$ from the standard BBN predictions. While this is slightly smaller than earlier estimates 2020ApJ ; Aver:2015iza ; Izotov:2014fga , inclusion of the primordial deuterium constraints lead to a $>2\sigma$ tension between the predicted number of neutrino species $N_{\rm eff}=2.41^{+0.19}_{-0.21}$ and the SM predicted value $N_{\rm eff}=3.045$, referred to as the Helium anomaly Matsumoto:2022tlr . With a large neutrino asymmetry in electron flavour $\xi_{e}=0.05^{+0.03}_{-0.03}$, it is however possible to obtain a large $N_{\rm eff}=3.22^{+0.33}_{-0.30}$ consistent with SM prediction within $1\sigma$ Matsumoto:2022tlr . A combined analysis of BBN and CMB observations have also found evidence for a large neutrino asymmetry in the early Universe at $\sim 2\sigma$ confidence level Burns:2022hkq .
@@ V Results and Discussion
$m_{\chi}^{2}(T)=m_{\chi}^{2}+\frac{g_{D}^{2}}{8}T^{2},$ (11)
with the correction being solely due to its gauge coupling. The top left and top right panel plots of Fig. 2 correspond to the benchmark points BP1, BP6 given in table 1 and table 2 respectively. Clearly, there exists a finite window $T_{n}<T<T_{s}$ during which $m_{\chi}>m_{\Phi_{2}}$ allowing the decay $\chi\rightarrow\Phi_{2}\nu$ and its conjugate process. We also show the sphaleron decoupling temperature $T_{\rm sph}\sim 130$ GeV to indicate that the large neutrino asymmetry is produced only at $T<T_{\rm sph}$ and hence can not get converted into baryon asymmetry. The bottom panel plots of the same figure shows the corresponding evolution of comoving number densities of $\chi,\bar{\chi},\Phi_{2}$ and $\Delta L_{\nu}$. Starting with an initial asymmetry $Y_{\chi}-Y_{\bar{\chi}}\neq 0$, the plots clearly depict the transfer of asymmetry in $\chi$ into a neutrino asymmetry during $T_{n}<T<T_{s}$. For $T<T_{n}$, the neutrino asymmetry saturates while $\chi,\bar{\chi}$ abundances continue to fall due to efficient annihilation $\chi\bar{\chi}\rightarrow Z_{D}Z_{D}$. Due to the finite temperature window and choice of Yukawa coupling $y_{1}$ controlling the decay rate, all dark asymmetry does not get transferred to neutrinos, leaving a remnant asymmetry. The remnant asymmetric relic for the chosen benchmark points agree with the observed DM relic. Note that, $\Phi_{2}$ decay for $T<T_{n}$ can not alter asymmetries in dark sector or neutrinos assuming that $\Phi_{2}$ does not store any asymmetry due to efficient $\Phi_{2}\rightarrow\Phi^{\dagger}_{2}$ interactions.
Fig. 3 shows the GW spectra for the benchmark points given in table 1 and 2. The same tables also contain the details of the GW related parameters used for calculating the spectra. Clearly, the peak frequencies can remain within the sensitivity of planned future experiment like LISA 2017arXiv170200786A , $\mu$ARES Sesana:2019vho , THEIA Garcia-Bellido:2021zgu , GAIA Garcia-Bellido:2021zgu , SKA Weltman:2018zrl and NANOGrav. The black colored violin-shaped points correspond to the recent NANOGrav data NANOGrav:2023gor . Due to the restrictions of having the FOPT below the electroweak scale, the peak frequencies do not come within reach of experiments like DECIGO Kawamura:2006up , BBO Yagi:2011wg among others.
@@
with $\tau$ being the observation time for a particular detector, which we consider to be 1 yr. We set the criteria SNR $>5(>1)$ for future (present) GW experiments to identify the parameter space. Clearly, the parameter space remains within reach of several GW experiments while at the same time offering complementary probe at future CMB experiments. As expected, lower values of DM mass requires the FOPT to occur at a lower scale to generate the finite temperature window for DM decay, pushing the corresponding peak frequencies towards the PTA ballpark. For the summary plot on the right panel of Fig. 4 and Fig. 5, $v_{D}$ is varied from $10^{-2}$ GeV to $10^{2}$ GeV, $\lambda_{\Phi_{1}\Phi_{2}}$ is varied from 1 to 3 and $Y_{\Delta\chi}$ is varied between $5\times 10^{-4}$ and $5\times 10^{-2}$, while $\lambda_{1}=\lambda_{2}=0.01$, $y_{1}$=9×$\times 10^{-9}$ and $g_{D}=0.4$. In Fig. 5, we show the zero temperature masses of dark sector particles namely DM and $\Phi_{2}$ satisfying all requirements and being within sensitivity of different GW experiments indicated in terms of colour code and point shape.
Figure 2: Top panel: Finite temperature masses of $\chi$ and $\Phi_{2}$ for BP1 (left) and BP6 (right) mentioned in table 1, 2 with $m_{\chi}$=18×$\times 10^{-3}$ GeV (left) and $m_{\chi}$=80 GeV (right). Bottom panel: Evolution of comoving number densities for BP1 (left) and BP6 (right) with $y_{1}=7.97\times 10^{-10}$ (left) and $y_{1}=1.13\times 10^{-7}$ (right).
$v_{D}$ $T_{c}$ $v_{c}$ $T_{n}$ $\mu_{\Phi_{2}}$ $m_{\Phi_{2}}$ $g_{D}$ $\lambda_{\Phi_{1}\Phi_{2}}$ $\lambda_{1}$ $\lambda_{2}$ $\alpha_{*}$ $\beta/\mathcal{H}_{*}$ $v_{J}$ $T_{\rm RH}$
(MeV) (MeV) (MeV) (MeV) (MeV) (MeV) (MeV)
BP1 20 6.13 18.40 2.57 5 23.30 0.4 2.59 0.01 0.01 1.24 56.06 0.94 2.68
BP2 40 12.27 36.80 5.10 10 45.79 0.4 2.59 0.01 0.01 1.29 51.65 0.94 5.38
BP3 60 17.98 56.40 7.22 20 68.57 0.6 2.39 0.01 0.01 1.65 44.10 0.95 8.12
Table 1: Benchmark points BP1, BP2, BP3 and other details involved in the GW spectrum calculations.
$v_{D}$ $T_{c}$ $v_{c}$ $T_{n}$ $\mu_{\Phi_{2}}$ $m_{\Phi_{2}}$ $g_{D}$ $\lambda_{\Phi_{1}\Phi_{2}}$ $\lambda_{1}$ $\lambda_{2}$ $\alpha_{*}$ $\beta/\mathcal{H}_{*}$ $v_{J}$ $T_{\rm RH}$
(GeV) (GeV) (GeV) (GeV) (GeV) (GeV) (GeV)
BP4 1 0.29 0.93 0.11 0.2 1.11 0.6 2.40 0.01 0.01 1.09 62.60 0.93 0.12
BP5 10 2.94 9.30 1.04 0.2 10.29 0.6 2.12 0.01 0.01 0.43 45.62 0.88 1.04
BP6 100 29.70 93.00 11.14 20 107.23 0.6 2.22 0.01 0.01 0.29 50.61 0.85 11.14
Table 2: Benchmark points BP4, BP5, BP6 and other details involved in the GW spectrum calculations.
Figure 3: GW spectrum corresponding to BP1 to BP3 (left) and BP1 to BP6 (right) of table 1, 2.

### v2

@@ IV Large neutrino asymmetry and $N_{\rm eff}$
In the SM, such neutrino asymmetries do not arise and hence there is no additional contribution to $N_{\rm eff}$. The sole contribution to $N_{\rm eff}$ in the SM comes from relativistic nature of light neutrinos. If neutrinos decouple instantaneously, then $N_{\rm eff}=3$. Considering non-instantaneous decoupling of SM neutrinos along with flavour oscillations and plasma correction of quantum electrodynamics, the SM $N_{\rm eff}$ value shifts to $N_{\rm eff}^{\rm SM}=3.045$ Mangano:2005cc ; Grohs:2015tfy ; deSalas:2016ztq  [^4] A few recent works Froustey:2020mcq ; Bennett:2020zkv ; Drewes:2024wbw obtained a slightly different prediction $N_{\rm eff}^{\rm SM}=3.044$.. A deviation from $N_{\rm eff}^{\rm SM}$ indicates presence of BSM physics either in terms of additional light degrees of freedom, new interactions of SM neutrinos or large neutrino asymmetry mentioned above. The current bound on $N_{\rm eff}$ from Planck 2018 data is given as $N_{\rm eff}=2.99^{+0.34}_{-0.33}$ at $2\sigma$ CL including baryon acoustic oscillation (BAO) data. This corresponds to $\Delta N_{\rm eff}\lesssim 0.285$. The latest DESI 2024 data give a slightly weaker bound $\Delta N_{\rm eff}\lesssim 0.4$ at $2\sigma$ CL DESI:2024mwx . Similar bound also exists from BBN considerations $2.3<{\rm N}_{\rm eff}<3.4$ at $95\%$ CL Cyburt:2015mya . Future CMB experiment CMB Stage IV (CMB-S4) is expected to reach a much better sensitivity of $\Delta{\rm N}_{\rm eff}={\rm N}_{\rm eff}-{\rm N}^{\rm SM}_{\rm eff}=0.06$ Abazajian:2019eic , taking it closer to the SM prediction. Another future experiment CMB-HD CMB-HD:2022bsz can probe $\Delta N_{\rm eff}$ upto $0.014$ at $1\sigma$.
In addition to enhancing $N_{\rm eff}$, a large neutrino asymmetry can also affect the BBN predictions of light nuclei abundance, as pointed out recently in the light of anomalous observations related to primordial Helium-4 ($^{4}{\rm He}$) abundance. The recent observations made by the Subaru Survey Matsumoto:2022tlr , together with previous observations, have indicated a shift in primordial $^{4}{\rm He}$ abundance $Y_{P}=0.2379^{+0.0031}_{-0.0030}$ from the standard BBN predictions. While this is slightly smaller than earlier estimates 2020ApJ ; Aver:2015iza ; Izotov:2014fga , inclusion of the primordial deuterium constraints lead to a $2\sigma$ tension between the predicted number of neutrino species $N_{\rm eff}=2.41^{+0.19}_{-0.21}$ and the SM predicted value $N_{\rm eff}=3.045$, referred to as the Helium anomaly Matsumoto:2022tlr . With a large neutrino asymmetry in electron flavour $\xi_{e}=0.05^{+0.03}_{-0.03}$, it is however possible to obtain a large $N_{\rm eff}=3.22^{+0.33}_{-0.30}$ consistent with SM prediction within $1\sigma$ Matsumoto:2022tlr . A combined analysis of BBN and CMB observations have also found evidence for a large neutrino asymmetry in the early Universe at $\sim 2\sigma$ confidence level Burns:2022hkq .
@@ V Results and Discussion
$m_{\chi}^{2}(T)=m_{\chi}^{2}+\frac{y^{2}_{\chi\Phi_{1}}}{16}T^{2},$ (11)
with the correction being solely due to its Yukawa coupling. The top left and top right panel plots of Fig. 2 correspond to the benchmark points BP1, BP6 given in table 1 and table 3 respectively. Due to the small values of Yukawa coupling $y_{\chi\Phi_{1}}$ chosen, $\chi$ does not receive noticeable thermal corrections to its mass. Clearly, there exists a finite window $T_{n}<T<T_{s}$ during which $m_{\chi}>m_{\Phi_{2}}$ allowing the decay $\chi\rightarrow\Phi_{2}\nu$ and its conjugate process. We also show the sphaleron decoupling temperature $T_{\rm sph}\sim 130$ GeV to indicate that the large neutrino asymmetry is produced only at $T<T_{\rm sph}$ and hence can not get converted into baryon asymmetry. The bottom panel plots of the same figure shows the corresponding evolution of comoving number densities of $\chi,\bar{\chi},\Phi_{2}$, $\Phi^{\dagger}_{2}$ and $\Delta L_{\nu}$. Starting with an initial asymmetry $Y_{\chi}-Y_{\bar{\chi}}\neq 0$, the plots clearly depict the transfer of asymmetry in $\chi$ into a neutrino asymmetry during $T_{n}<T<T_{s}$. For $T<T_{n}$, the neutrino asymmetry saturates while $\chi,\bar{\chi}$ abundances continue to fall due to efficient annihilation $\chi\bar{\chi}\rightarrow\Phi_{1}\Phi_{1}$. Due to the finite temperature window and choice of Yukawa coupling $y_{1}$ controlling the decay rate, all dark asymmetry does not get transferred to neutrinos, leaving a remnant asymmetry. The remnant asymmetric relic for the chosen benchmark points agree with the observed DM relic. Note that, $\Phi_{2}$ decay for $T<T_{n}$ can not alter asymmetries in dark sector or neutrinos as $\Phi_{2}$ does not store any asymmetry due to efficient $\Phi_{2}\rightarrow\Phi^{\dagger}_{2}$ interactions. This is seen from the evolutions of $\Phi_{2}$ and $\Phi^{\dagger}_{2}$ which almost coincide with each other. The details of this conversion process can be found in Appendix C.
[…]
@@
$\rho=\sqrt{\tau\,\int_{f_{\text{min}}}^{f_{\text{max}}}\,df\,\left[\frac{\Omega_{\text{GW}}(f)\,h^{2}}{\Omega_{\text{expt}}(f)\,h^{2}}\right]^{2}}\,$, (12)
@@
[…]
$v_{D}$ $\mu_{\Phi_{2}}$ $m_{\Phi_{2}}$ $\lambda_{\Phi_{1}\Phi_{2}}$ $\lambda_{1}$ $\lambda_{2}$ $\lambda_{\chi\Phi_{1}}$ $\lambda_{12}$
(MeV) (MeV) (MeV)
BP1 20 5 20.25 1.92 0.01 0.01 0.01 $10^{-3}$
BP2 40 10 40.49 1.92 0.01 0.01 0.01 $10^{-3}$
BP3 60 20 63.48 2.01 0.01 0.01 0.01 $10^{-3}$
Table 1: Benchmark points BP1, BP2, BP3 for different input model parameters.
$T_{c}$ $v_{c}$ $T_{n}$ $\alpha_{*}$ $\beta/\mathcal{H}_{*}$ $v_{J}$ $T_{\rm RH}$
(MeV) (MeV) (MeV) (MeV)
BP1 6.45 19.00 2.70 0.12 40.37 0.79 2.70
BP2 12.92 38.00 5.45 0.12 46.41 0.79 5.45
BP3 19.87 57.00 9.04 0.08 56.93 0.76 9.04
Table 2: Benchmark points BP1, BP2, BP3 with other details involved in the GW spectrum calculations.
$v_{D}$ $\mu_{\Phi_{2}}$ $m_{\Phi_{2}}$ $\lambda_{\Phi_{1}\Phi_{2}}$ $\lambda_{1}$ $\lambda_{2}$ $\lambda_{\chi\Phi_{1}}$ $\lambda_{12}$
(GeV) (GeV) (GeV)
BP4 1 0.2 0.98 1.86 0.01 0.01 0.01 $10^{-3}$
BP5 10 0.2 9.34 1.74 0.01 0.01 0.01 $10^{-3}$
BP6 100 20 98.36 1.86 0.01 0.01 0.01 $10^{-3}$
Table 3: Benchmark points BP4, BP5, BP6 for different input model parameters.
$T_{c}$ $v_{c}$ $T_{n}$ $\alpha_{*}$ $\beta/\mathcal{H}_{*}$ $v_{J}$ $T_{\rm RH}$
(GeV) (GeV) (GeV) (GeV)
BP4 0.32 0.95 0.13 0.12 63.64 0.78 0.13
BP5 3.14 9.50 1.12 0.23 51.01 0.83 1.12
BP6 32.14 95.00 12.30 0.15 33.64 0.80 12.30
Table 4: Benchmark points BP4, BP5, BP6 with other details involved in the GW spectrum calculations.
Figure 3: GW spectra for BP1 to BP6 of table 2, 4.